OpenStudy (anonymous):

Is \((AB)^{-1}=A^{-1}B^{-1}\) for square matrices A and B?

OpenStudy (turingtest):

yes

OpenStudy (anonymous):

really? I couldn't find anything that says so. Wikipedia or google...

jimthompson5910 (jim_thompson5910):

Shouldn't it be \(\Large (AB)^{-1}=B^{-1}A^{-1}\) ???

OpenStudy (anonymous):

Yeah that's what I thought too.

OpenStudy (anonymous):

I guess the matrices I tried it out on were trivial...

jimthompson5910 (jim_thompson5910):

AB = C (AB)^(-1)(AB) = (AB)^(-1)C I = (AB)^(-1)AB I = B^(-1)A^(-1)AB I = B^(-1)[A^(-1)A]B I = B^(-1)[I]B I = B^(-1)B I = I It's explained better here http://mathworld.wolfram.com/MatrixInverse.html

OpenStudy (turingtest):

Oh I totally didn't even read it... sorry :/